OpenStudy (anonymous):

Algebra help please!!<333 last question and ill be done. I dont get it! :/ The volume of gas varies directly with temperature and inversely with pressure. Volume is 100 m^3 when temperature is 1508 degrees. and pressure is 15 lb/cm^2 what is the volume when the temperature is 2508 degrees and the pressure is 20 lb/cm^2 ?

OpenStudy (zehanz):

You have to convert the sentences to a formula:\[V=c \frac{ T }{ p }\] Put in the numbers of the first situation to calculate the constant c. Then put in the numbers of the second situation. You can calculate V now.

OpenStudy (anonymous):

:D thanks

OpenStudy (zehanz):

Hope you understand now...

OpenStudy (anonymous):

kinda. Well, i do have a question. the constant C part. im kinda stuck o.o

OpenStudy (anonymous):

im kinda just stuck at how to plug in this first situation. from there ill be ok.

OpenStudy (zehanz):

First situation:\[100=c \frac{ 1508 }{ 15 }\]You can get c by multiplying both sides with 15 and divide by 1508.

OpenStudy (anonymous):

thanks. alright, let me try this.

OpenStudy (anonymous):

by both sides you mean the 100 too right? Sorry. I ask so many questions lol D:

OpenStudy (zehanz):

Yes!

OpenStudy (anonymous):

ok well, i did that. Now...i dont divide that part by the 1508 right? :/ like i multiplied the 100 by 15 and got 1500

OpenStudy (zehanz):

c=100*15/1508=1500/1508 or about 0.9947

OpenStudy (anonymous):

yes, i did that. so thats not wrong? I get nervous about math and i thought i did it wrong o.o ah good, so then, from there...

OpenStudy (zehanz):

So now try putting the new numbers in!

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

i put the new numbers in the formula right?

OpenStudy (zehanz):

Uhuh

OpenStudy (anonymous):

do i do anything with the number i just got. and the new number i got was the constant c right? would i just plug it in like, v=2508/20 or do i put the c in too.